Derive the consumer satisfaction level in the law of equal marginal utility when there are two commodities A&B the price of A=2 and B=3 Rs the consumer income is 24.
1
Expert's answer
2020-10-19T13:37:59-0400
"2A+3B=24"

"\frac {MU_A}{2}=\frac {MU_B}{3}"